OpenStudy (pottersheep):

Can someone show me how to find the Value of "a" in this rather simple arithmetic means problem? Grade 11... Totally blanked out as a result of the break :P

OpenStudy (pottersheep):

Determine the missing terms in the arithmetic sequence which has T1 = -a/2 T6 = 2a

OpenStudy (pottersheep):

I deleted the old one by accident my bad...

OpenStudy (pottersheep):

for a I got -a/2 d I got : a/2

OpenStudy (pottersheep):

not sure if i am correct..

OpenStudy (amistre64):

t1 = -a/2 t2=-a/2+d t3=-a/2+2d t4=-a/2+3d t5=-a/2+4d t6=-a/2+5d = 2a is what im thinking of

OpenStudy (pottersheep):

Yup amistre64 that's what I based it on as well

OpenStudy (anonymous):

The elements of an arithmetic sequence are separated by a common difference. If you can find that difference, then you can calculate the value of any element in the sequence. The first element of the sequence is \(\frac{-a}{2}\) and for any arithmetiv sequence with common difference d, \[a_n = a_{n-1} + d(n-1)\]

OpenStudy (pottersheep):

then I did t6 = a + 5d so 2a - ( - a/2) = 5d d = a/2

OpenStudy (anonymous):

right

OpenStudy (amistre64):

5d = 2a + a/2 = (4a+a)/2 = 5a/2 d = 5a/2.5 d = a/2 i agree so far lol

OpenStudy (anonymous):

i did a little mistake in my rule though, instead of \(a_{n-1}\)it should be \(a_1\)

OpenStudy (pottersheep):

True

OpenStudy (anonymous):

i think your answer is right

OpenStudy (pottersheep):

It's right so far I guess, the final answers given in my book are: 0, 0.75, 1.1 , 1.45 as the missing terms

OpenStudy (pottersheep):

Oh well thanks for helping ...

OpenStudy (pottersheep):

I'll think a little more and see if I get anywhere:P

OpenStudy (pottersheep):

GOT IT :D